Robert invests $8000 in a bank. Part is invested at 7% simple interest and part at 9% simple interest. In one year Robert earns $660 in interest. How much did he invest at each amount?
Answer by checkley77(12569)

.09X+.07(8,000-X)=660
.09X+560-.07X=660
.02X=660-560
.02X=100
X=100/.02
X=5,000 AMOUNT INVESTED @ 9%.
8,000-5,000=3,000 AMOUNT INVESTED @ 7%.
PROOF:
.09*5,000+.07*3,000=660
450+210=660
660=660